ʃ1/(2/3x^5+2x-1 dx
1
Expert's answer
2019-04-03T10:03:56-0400
"\int({2 \over 3}x^5+2x-1)dx={2 \over 3}({x^6 \over 6})+2({x^2 \over 2})-x+C="

"={1 \over 9}x^6+x^2-x+C="